How many grams of sodium hydroxide are required to neutralize 2500.0 mL of 5.493 M of sulfuric acid? A balance equation maybe required.
1
Expert's answer
2020-05-05T03:36:54-0400

"2NaOH +H_2SO_4 \to Na_2SO_4+2H_2O"


Thus, 2 moles Sodium hydroxide reacts with 1 mole of Sulfuric acid.

Let x moles react with 2500.0 mL of 5.493 M of sulfuric acid

"\implies x=2.5*5.493/2=6.86625 moles"

Molar mass of NaOH is 40 grams.

Thus, "40*6.86625\n=274.65 gms." are needed.